### v2

@@ Abstract
${}^{\dagger}$orresponding to Jifeng Dai ¡daijifeng@tsinghua.edu.cn¿.
@@ 4 The All-Seeing Model v2
ASMv2 is a powerful Multi-modal Large Language Model (MLLM), which integrates the Relation Conversation (ReC) ability while maintaining powerful general capabilities. Specifically, it follows the model architecture of LLaVA-1.5 [49], comprising a vision encoder, a vision-language connector, and a language model. This model can deal with three types of relation tasks, including (1) Relation Conversation, which requires the model to link all objects and predicates mentioned in the response to the corresponding regions in the image. (2) Open-ended Scene Graph Generation, which requires the model to generate a scene graph based on the given image in an open-ended manner; (3) Predicate Classification, which requires the model to generate a scene graph given the ground-truth object labels and localization; In addition, our model is also capable of multi-modality dialogue tasks such as Image Captioning, Visual Question Answering, and Multi-turn conversation. Since the ReC task requires the model to link the objects and predicates to the corresponding regions in the image, our ASMv2 is also endowed with grounding and referring capabilities and exhibits state-of-the-art performance on region-level tasks, including Referring Expression Comprehension, Region Captioning, and Referring Question Answering.
@@ 4.2 Model Training
The training process of ASMv2 is divided into two stages, with each stage comprising a pre-training phase and an instruction-tuning phase. The first stage is designed to enable the model to effectively understand visual information at the image level. The pre-training phrase utilizes 595K samples from CC3M [69] filtered by LLaVA [50] while the instruction-tuning phrase utilizes a blend of 665K samples from LLaVA-Mix [49]. We update the data format of the region-level data [30, 57, 31] in LLaVA-Mix to the format introduced in Sec. 4.1. The second stage trains the model with a mixture of image-level data and region-level data, which enables the model to comprehend the visual information at the region level, facilitating effective grounding of objects and predicates within sentences. The pre-training phrase employs 5M samples from CC12M [69] filtered by BLIP [36], 10M filtered samples from AS-1B [80], and 15M filtered samples from GRiT [62]. The instruction-tuning phase employs 4M samples collected from a variety of sources, including image-level datasets [50, 49, 99, 9, 48, 47, 29, 63, 56, 12, 4], region-level datasets [30, 57, 31, 80, 91, 100] and our proposed AS-V2 dataset. The summary of these datasets is presented in Tab. 9.
@@ Appendix 0.B The All-Seeing Dataset v2
messages = [{"role":"system", "content": f"""You are an AI visual assistant that can analyze a single image. You receive one image and five sentences, each describing this image you are observing. In addition, specific object locations within the image are given, along with detailed coordinates. These coordinates are in the form of bounding boxes, represented as [x1, y1, x2, y2] with int numbers ranging from 0 to 999. These values correspond to the top left x, top left y, bottom right x, and bottom right y. Note that these coordinates are normalized. Besides, the scene graph of this image is also provided as a list of tuples. Each tuple is represented as (subject, bounding box of the subject, object, bounding box of the object, predicate).
Using the provided caption, bounding box, and scene graph information, describe the scene in a detailed manner. If there are errors in the caption, please ignore them and do not point them out in your description.
Instead of directly mentioning the bounding box coordinates, utilize this data to explain the scene using natural language with its bounding box in the format like "<ref>object</ref><box>[[x1, y1, x2, y2]]</box>". When mentioning the predicate between two objects, you should mention it in the format like "<pred>predicate</pred><box>[[x1, y1, x2, y2]]</box><box>[[x3, y3, x4, y4]]</box>", where "<box>[[x1, y1, x2, y2]]</box>" denotes the bounding box coordinates of the subject and "<box>[[x3, y3, x4, y4]]</box>" denotes the bounding box coordinates of the object. Include details like object counts, position of the objects, relative position between the objects.
When using the information from the caption, coordinates, or scene graph, directly explain the scene, and do not mention that the information source is the caption or the bounding box or the scene graph. You should mention all tuples and predicates included in the scene graph in the generated caption. Make sure that the box following the <pred>predicate</pred> has already been mentioned after a <ref>object</ref>."""}
for sample in fewshot_samples:
@@
$\cdots$
Response type 3: Conversation

### v3

@@ Abstract
Corresponding to Jifeng Dai ¡daijifeng@tsinghua.edu.cn¿.
@@ 4 The All-Seeing Model v2
ASMv2 is a powerful Multi-modal Large Language Model (MLLM), which integrates the Relation Conversation (ReC) ability while maintaining powerful general capabilities. Specifically, it follows the model architecture of LLaVA-1.5 [50], comprising a vision encoder, a vision-language connector, and a language model. The Vicuna-13B [103] and CLIP-ViT-L-336px [20] are utilized as the language and vision backbone. This model can deal with three types of relation tasks, including (1) Relation Conversation, which requires the model to link all objects and predicates mentioned in the response to the corresponding regions in the image. (2) Open-ended Scene Graph Generation, which requires the model to generate a scene graph based on the given image in an open-ended manner; (3) Predicate Classification, which requires the model to generate a scene graph given the ground-truth object labels and localization; In addition, our model is also capable of multi-modality dialogue tasks such as Image Captioning, Visual Question Answering, and Multi-turn conversation. Since the ReC task requires the model to link the objects and predicates to the corresponding regions in the image, our ASMv2 is also endowed with grounding and referring capabilities and exhibits state-of-the-art performance on region-level tasks.
@@ 4.2 Model Training
The training process of ASMv2 is divided into two stages, with each stage comprising a pre-training phase and an instruction-tuning phase. The first stage is designed to enable the model to effectively understand visual information at the image level. The pre-training phrase utilizes 595K samples from CC3M [70] filtered by LLaVA [51] while the instruction-tuning phrase utilizes a blend of 665K samples from LLaVA-1.5 [50]. We update the data format of the region-level data [31, 58, 32] in LLaVA-1.5 to the format introduced in Sec. 4.1. The second stage trains the model with a mixture of image-level data and region-level data, which enables the model to comprehend the visual information at the region level, facilitating effective grounding of objects and predicates within sentences. The pre-training phrase employs 5M samples from CC12M [70] filtered by BLIP [37], 10M filtered samples from AS-1B [81], and 15M filtered samples from GRiT [63]. The instruction-tuning phase employs 4M samples collected from a variety of sources, including image-level datasets [51, 50, 100, 9, 49, 48, 30, 64, 57, 12, 4], region-level datasets [31, 58, 32, 81, 92, 101] and our proposed AS-V2 dataset. The summary of these datasets is presented in Tab. 9.
@@ Appendix 0.B The All-Seeing Dataset v2
messages = [{"role":"system", "content": f"""You are an AI visual assistant that can analyze a single image. You receive one image and five sentences, each describing this image you are observing. In addition, specific object locations within the image are given, along with detailed coordinates. These coordinates are in the form of bounding boxes, represented as [x1, y1, x2, y2] with int numbers ranging from 0 to 999. These values correspond to the top left x, top left y, bottom right x, and bottom right y. Note that these coordinates are normalized. Besides, the scene graph of this image is also provided as a list of tuples. Each tuple is represented as (subject, bounding box of the subject, object, bounding box of the object, predicate). Using the provided caption, bounding box, and scene graph information, describe the scene in a detailed manner. If there are errors in the caption, please ignore them and do not point them out in your description. Instead of directly mentioning the bounding box coordinates, utilize this data to explain the scene using natural language with its bounding box in the format like "<ref>object</ref><box>[[x1, y1, x2, y2]]</box>". When mentioning the predicate between two objects, you should mention it in the format like "<pred>predicate</pred><box>[[x1, y1, x2, y2]]</box><box>[[x3, y3, x4, y4]]</box>", where "<box>[[x1, y1, x2, y2]]</box>" denotes the bounding box coordinates of the subject and "<box>[[x3, y3, x4, y4]]</box>" denotes the bounding box coordinates of the object. Include details like object counts, position of the objects, relative position between the objects. When using the information from the caption, coordinates, or scene graph, directly explain the scene, and do not mention that the information source is the caption or the bounding box or the scene graph. You should mention all tuples and predicates included in the scene graph in the generated caption. Make sure that the box following the <pred>predicate</pred> has already been mentioned after a <ref>object</ref>."""} ] for sample in fewshot_samples:
@@
$\cdots$ Response type 3: Conversation